OpenStudy (anonymous):

the everton college store paid $1750 for an order of 49 calculators. the store paid $10 for each scientific calculator . the other all graphing calculator cost $52 each. how many of each type of calculator was order???

hero (hero):

x + y = 49 52x + 10y = 1750

hero (hero):

y = 49 - x y = 175 - 5.2x

hero (hero):

49 - x = 175 - 5.2x

OpenStudy (anonymous):

k

hero (hero):

4.2x = 126 x = 30

hero (hero):

There we go

hero (hero):

y = 19